Suppose x has the pmf f(x)={kx for x=1,2,3,4

{0,elsewhere

Find the value of the constant k hence obtain the mean and the variance of x


1
Expert's answer
2021-07-13T15:16:18-0400
"k(1)+k(2)+k(3)+k(4)=10k=1"

"k=0.1"

"mean=E[X]=1(0.1)+2(0.2)+3(0.3)+4(0.4)=3"

"E[X^2]=1^2(0.1)+2^2(0.2)+3^2(0.3)+4^2(0.4)=10"

"V(X)=E[X^2]-(E[X])^2=10-(3)^2=1"